# A recursive approach for the determination of the nice sections of width three having a 4-crown stack as retract

### v1

@@ 5 Application
  • P = 1111: Criterion 5.1.1 prevents $P(4)$ from being a candidate for an $s$-base because the 6-crown $P(0,1)$ does not have a suitable retract. Furthermore, selecting $P(1\rightarrow 4)$ as $s$-base does not work due to Criterion 5.1.5, because $P(3,4)\simeq 1$ is marked with “n” in Table 1. $1111$ does thus have no retractive down-split fulfilling (5), and because the poset is self-dual, we conclude that it does not have a 4-crown stack as retract.
  • $P=1101$: Again, $P(1\rightarrow 4)$ fails as an $s$-base due to Criterion 5.1.5. $1101$ thus does not have a retractive down-split fulfilling (5).
  • $P=1011$: $P(1\rightarrow 4)$, $P(3,4)$, or $P(4)$ being a successful $s$-base is prevented by Criterion 5.1.4 together with Lemma 5.2, Criterion 5.1.3, and Criterion 5.1.1, respectively. Also $1011$ does thus not have a retractive down-split fulfilling (5). Therefore, neither $1101$ nor $1011$ has a 4-crown stack as retract.
@@
  • $P=11111$: Due to the Criteria 5.1.5 and 5.1.3, the upper segments $P(1\rightarrow 5)$ and $P(4,5)$ cannot work as $s$-bases. The self-dual poset has thus no 4-crown stack as retract.
@@
  • $P=11011$: Criterion 5.1.5 prevents $P(1\rightarrow 5)$ from being a successful $s$-base. Because $11011$ is self-dual, the poset cannot have a 4-crown stack as retract.
  • $P=10111$: Figure 2 shows that the 6-crown stack $Q:P(2\rightarrow 5)=111$ has a retraction $t:Q\rightarrow T$ with $T$ being a 4-crown stack and $t^{-1}(a)=\{a\}$ for a point $a\in T(0)$. With $s:P(0)\rightarrow P(0)$ being a mapping with $s[P(0)]$ being a 2-antichain, apply Corollary 4.6 with $k=1$.
@@
  • $P=101001,100001$: Let $P$ be one of these posets. With $Q:=P(3\rightarrow 6)=001$, select a point $d\in Q(0)$ and let $s$ be the retraction of $Q\setminus\{d\}$ onto a 4-crown $S$ shown in Figure 6 on the right with $a\in S\cap Q(0)$ as in the figure. Furthermore, let $t$ be the retraction of $P(0\rightarrow 2)=10$ onto a 4-crown $T$ as shown in Figure 5a with $v\in T\cap P(2)$ as in the figure. Applying Theorem 4.2 on $P(0\rightarrow 2)$, we can assume $v\lessdot a$ and $v\not<d$ and get a retractive down-split $(2,\{d\},s,t)$ fulfilling (5).
@@
The remaining posets of height six are $111011$, $111001$, $101011$, $110001$ and their duals and the self-dual poset $110011$. They do not have a 4-crown stack as retract:
@@
None of the posets $111011$, $111001$, $101011$, $110001$, and $110011$ has a 4-crown stack as retract.
@@
Now let $P$ be any of the five posets and $r$ a retraction onto a 4-crown stack $R$. With $S:=P(3\rightarrow 6)\cap R$, we show in a first step that $S$ cannot be a 4-crown stack.
$P(0\rightarrow 3)=111,110$: $S$ being a 4-crown stack yields a point $x\in P(3)$ with $v:=r(x)\in P(0\rightarrow 2)=11$. Let $\{v,w\}$ be the level set of $R$ containing $v$. $x\in P(3)$ yields $r[P(0,1)]\leq v$, hence $w\notin P(0,1)$. But then, $r(z)<\{v,w\}$ holds for at least five of the points $z\in P(0,1)$, and $P(0,1)$ cannot be mapped onto a 2-antichain.
$P=101011$: $S$ being a 4-crown stack requires due to Lemma 5.2 two points $x,y\in P(3)$ to be mapped into $P(0\rightarrow 2)=10$. The points $r(x)$ and $r(y)$ belong thus to $R(0,1)$. Three cases are possible, each of it leading to a contradiction:
@@
$S$ is thus not a 4-crown stack. This requires a level set $R(\ell)=\{a,b\}$ with $a\in P(2)$ and $b\in P(3)$. Clearly, $\ell=1$. Due to $P(1)<b$, no point of $P(1)$ can be mapped to $a$. And if a point of $P(1)$ is mapped to $b$, implication (4) yields $R(2\rightarrow h_{R})$ being a retract of $P(4\rightarrow 6)=11$ which is impossible. But if no point of $P(1)$ can be mapped to $R(1)$, then $r[P(0,1)]=R(0)$ which is impossible, too.

### v2

@@ 5 Application
  • P = 1111: Criterion 5.1.5 prevents $P(4)$ from being a candidate for an $s$-base because the 6-crown $P(0,1)$ does not have a suitable retract. Furthermore, selecting $P(1\rightarrow 4)$ as $s$-base does not work due to Criterion 5.1.2, because $P(3,4)\simeq 1$ is marked with “n” in Table 1. $1111$ does thus have no retractive down-split fulfilling (3), and because the poset is self-dual, we conclude that it does not have a 4-crown stack as retract.
  • $P=1101$: Again, $P(1\rightarrow 4)$ fails as an $s$-base due to Criterion 5.1.2. Therefore, $1101$ does not have a retractive down-split fulfilling (3).
  • $P=1011$: $P(1\rightarrow 4)$, $P(3,4)$, or $P(4)$ being a successful $s$-base is prevented by Criterion 5.1.1 together with Lemma 5.2, Criterion 5.1.4, and Criterion 5.1.5, respectively. Also $1011$ does thus not have a retractive down-split fulfilling (3). Therefore, neither $1101$ nor $1011$ has a 4-crown stack as retract.
@@
  • $P=11111$: Due to the Criteria 5.1.2 and 5.1.4, the upper segments $P(1\rightarrow 5)$ and $P(4,5)$ cannot be work as $s$-bases. The self-dual poset has thus no 4-crown stack as retract.
@@
  • $P=11011$: Criterion 5.1.2 prevents $P(1\rightarrow 5)$ from being a successful $s$-base. Because $11011$ is self-dual, the poset cannot have a 4-crown stack as retract.
  • $P=10111$: Figure 2 shows that the 6-crown stack $Q:P(2\rightarrow 5)=111$ has a retraction $t:Q\rightarrow T$ with $T$ being a 4-crown stack and $t^{-1}(a)=\{a\}$ for a point $a\in T(0)$. With $s:P(0)\rightarrow P(0)$ being a mapping such that $s[P(0)]$ is a 2-antichain, apply Corollary 4.6 with $k=1$.
@@
  • $P=101001,100001$: Let $P$ be one of these posets. With $Q:=P(3\rightarrow 6)=001$, select a point $d\in Q(0)$ and let $s$ be the retraction of $Q\setminus\{d\}$ onto a 4-crown $S$ shown in Figure 8d with $a\in S\cap Q(0)$ as in the figure. Furthermore, let $t$ be the retraction of $P(0\rightarrow 2)=10$ onto a 4-crown $T$ as shown in Figure 7a with $v\in T\cap P(2)$ as in the figure. Applying Theorem 4.2 on $P(0\rightarrow 2)$, we can assume $v\lessdot a$ and $v\not<d$ and get a retractive down-split $(2,\{d\},s,t)$ fulfilling (3).
@@
For $P=110011$, the Criteria 5.1.2 and 5.1.5 prevent $P(1\rightarrow 6)$ and $P(6)$ from being successful $s$-bases; the self-dual poset has thus now 4-crown stack as retract. The remaining posets of height six are $111011$, $111001$, $101011$, $110001$, and their duals. They do not have a 4-crown stack as retract:
@@
None of the posets $111011$, $111001$, $101011$, and $110001$ has a 4-crown stack as retract.
@@
Now let $P$ be any of the four posets and $r$ a retraction onto a 4-crown stack $R$. With $S:=P(3\rightarrow 6)\cap R$, we show in a first step that $S$ cannot be a 4-crown stack.
$P(0\rightarrow 3)=111,110$: $P(3\rightarrow 6)$ does not have a 4-crown stack as retract. $S$ being a 4-crown stack requires thus a point $x\in P(3)$ with $v:=r(x)\in P(0\rightarrow 2)=11$. Let $\{v,w\}$ be the level set of $R$ containing $v$. $x\in P(3)$ yields $r[P(0,1)]\leq v$, hence $w\notin P(0,1)$. But then, $r(z)<\{v,w\}$ holds for at least five of the points $z\in P(0,1)$, and $P(0,1)$ cannot be mapped onto a 2-antichain.
$P=101011$: $S$ being a 4-crown stack requires due to Lemma 5.2 two points $x,y\in P(3)$ to be mapped into $P(0\rightarrow 2)=10$. The points $r(x)$ and $r(y)$ belong thus to $R(0,1)$. Three cases are possible, each leads to a contradiction:
@@
$S$ is thus not a 4-crown stack. This requires a level set $R(\ell)=\{a,b\}$ with $a\in P(2)$ and $b\in P(3)$. Clearly, $\ell=1$. Due to $P(1)<b$, no point of $P(1)$ can be mapped to $a$. And if a point of $P(1)$ is mapped to $b$, (2) yields $R(2\rightarrow h_{R})$ being a retract of $P(4\rightarrow 6)=11$ which is impossible. But if no point of $P(1)$ can be mapped to $R(1)$, then $r[P(0,1)]=R(0)$ which is impossible, too.